stis status n.
Skip this Video
Download Presentation
STIS Status

Loading in 2 Seconds...

play fullscreen
1 / 9

STIS Status - PowerPoint PPT Presentation

  • Uploaded on

STIS Status. TIPS July 16, 2009. STIS Status. STIS MIE anomaly What is known STIS CCD status before the anomaly STIS MAMA SMOV status. STIS MIE Anomaly. STIS currently suspended

I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
Download Presentation

PowerPoint Slideshow about 'STIS Status' - melosa

Download Now An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.

- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
stis status

STIS Status


July 16, 2009

stis status1
STIS Status
  • STIS MIE anomaly
    • What is known
  • STIS CCD status before the anomaly
  • STIS MAMA SMOV status
stis mie anomaly
STIS MIE Anomaly
  • STIS currently suspended
  • Original suspend (day 187): Main STIS CS computer unable to get timecode update from MIE (MAMA Interface Electronics) that was needed to populate the header of a CCD exposure.
    • Originally suspected that single event upset had hung up the MIE 486.
  • 2nd suspend (day 190): Suspended while attempting to copy MIE memory to CS. Done before full reset in attempt to preserve MIE memory
    • Theorized that problem with MIE dump was caused by lack of full reset, and hoped a full reset would clear the problem.
    • About 1/2 of MIE memory was successfully copied.
  • 3rd suspend (191): Did full reset and successfully recovered STIS to operate, but suspended while attempting to copy MIE memory to CS.
    • MIE copy to CS seems to have succeeded, but still suspended
    • Recovery included successful queries of MIE for time code update (I.e., the original problem from day 187 did not recur).
    • Observed events suggest transient failure on MIE FIFO Flags interface between CS and MIE?

STIS MIE Anomaly - cont

  • What we know:
    • The three anomalies occurred during three different circumstances and have three different and - in some aspects - inconsistent signatures
    • Observed events suggest transient failure on MIE FIFO Flags interface between CS and MIE
  • Investigation is continuing
    • Perhaps MIE dump problem is separate from original time code update problem?
    • STIS still suspended
  • Hope to allow resumption of CCD operations soon, while analysis continues.
    • See if original anomaly reoccurs.
    • CCD observations could, in principle work around MIE problems
    • Since anomalies are not yet understood, implications for MAMA observations are not yet known
stis ccd read noise
STIS CCD Read Noise
  • STIS CCD read noise slightly increased from ~ 5.4 e- in 2004 to ~ 5.7 e- now
  • Side 2 read-noise includes ~ 1 DN of semi-coherent 17 kHz pattern noise
    • This pattern noise appears to be unchanged after SM4
stis ccd dark current
STIS CCD Dark Current
  • CCD housing temperature (OCCDHT) surrogate for CCD chip T
  • For comparison these plots scale darks to housing temperature of 18 C
  • Dark rate increases by 7%/degree of housing temperature
  • 2009 aft-shroud hotter than 2004
    • When MAMAs on, OCCDHT 22-25 C
    • With MAMAs off, OCCDHT 16-18 C
  • RN between 0.009 and 0.013 e-/s, fluctuating with temperature
tds cte effects
TDS & CTE effects
  • Standard monitor target AGK+81o266 (*’s) shows large scatter at low flux levels (right panel), even in 2004 - correlates with T
  • Need to separate T dependence of CTE before deriving final throughput estimates
CTE extrapolated to zero background as a function of flux level
  • 2009 points are at colder T than older data
    • Expect reduced CTE losses at lower T
mama smov status
    • FUV MAMA HV ramped to full values
    • Fold test & flat look as expected
    • Dark current perhaps lower than expected
    • Initial steps of NUV MAMA HV recovery completed
    • Have not yet ramped NUV HV to level where tube generates real counts.
  • Still need to do initial internal lamp exposures and external exposures with both MAMAs
  • Once MAMAs cleared for operations, most SMOV activities for either or both could be completed in a week if time available.